# CoT-Self-Instruct: Building high-quality synthetic prompts for reasoning and non-reasoning tasks

### v1

@@ Abstract
We propose CoT-Self-Instruct, a synthetic data generation method that instructs LLMs to first reason and plan via Chain-of-Thought (CoT) based on the given seed tasks, and then to generate a new synthetic prompt of similar quality and complexity for use in LLM training, followed by filtering for high-quality data with automatic metrics. In verifiable reasoning, our synthetic data significantly outperforms existing training datasets, such as s1k and OpenMathReasoning, across MATH500, AMC23, AIME24 and GPQA-Diamond. For non-verifiable instruction-following tasks, our method surpasses the performance of human or standard self-instruct prompts on both AlpacaEval 2.0 and Arena-Hard.
@@ 1 Introduction
Synthetic data is artificially generated to replicate the characteristics and patterns of real-world data. One innovative approach to creating such data is the Self-Instruct method (Wang et al., 2022a), which utilizes LLMs themselves to generate instruction-following examples. This method begins by selecting a small set of seed instruction-following samples, which are then used to prompt LLMs to produce additional demonstrations in a similar format. Since then a number of variants have been introduced that increase the complexity of queries (Liu et al., 2023; Zeng et al., 2024), maintain semantic diversity (Ding et al., 2023), scale the synthetic data (Yuan et al., 2023), and use these methods in self-improvement loops (Yuan et al., 2024). However, a significant challenge with these approaches is to ensure the quality and effectiveness of the generated data for language model training. Overall, generating high-quality synthetic data and optimizing its use for both reasoning and non-reasoning tasks still remains insufficiently understood.
@@
In this paper, we present Chain-of-Thought(CoT)-Self-Instruct, a method that both (i) uses reasoning to help create high quality synthetic data; and (ii) self-filters the created data to only keep the highest quality ones, see Figure 1. We show the efficacy of this approach for creating both verifiable reasoning data and non-verifiable instruction following tasks, where in both cases using Chain-of-Thought (CoT) to help generate the examples outperforms those generated without CoT. To curate high quality verifiable data we introduce Answer-Consistency, which discards examples where the CoT-Self-Instruct-generated answer does not match the majority vote solution of the LLM, with the assumption that those examples are either incorrectly labeled or too difficult. For non-verifiable data we use the recent Rejecting Instruction Preferences (RIP) (Yu et al., 2025) method which measures the quality of prompts based on the distribution of reward model scores from LLM solutions. In both cases, filtering provides further gains. For reasoning tasks, CoT-Self-Instruct generated training data outperforms Self-Instruct and existing datasets such as s1k (Muennighoff et al., 2025) and OpenMathReasoning (Moshkov et al., 2025) across MATH500, AMC23, AIME24 and GPQA-Diamond. For non-reasoning tasks, it outperforms human data from WildChat (Zhao et al., 2024) and Self-Instruct, whether filtered or not, on both AlpacaEval 2 and ArenaHard. Overall, models trained with CoT-Self-Instruct generated training data provided the best results of the methods we tested.
@@
Figure 1: CoT-Self-Instruct. Our method first prompts LLMs to reason and generate new instructions given seed prompts, followed by automatically curating high-quality data using either Answer-Consistency for verifiable reasoning tasks, or RIP (Yu et al., 2025) for non-verifiable tasks.
@@ 2 Related Work
Synthetic data is produced using algorithms (Saxton et al., 2019), generative models (Borisov et al., 2022; Meng et al., 2022), or simulations (Vezhnevets et al., 2023), rather than being directly created by humans (Liu et al., 2024). It presents a promising solution for training models, particularly in scenarios where real-world data is scarce, expensive, or difficult to obtain. Self-Instruct (Wang et al., 2022a) first proposed a framework that lets a language model be prompted with seed data as few-shot examples in order to generate new synthetic data. Such data has been used to then self-train language models, e.g. in the Self-Rewarding framework (Yuan et al., 2024). Evol Instruct (Zeng et al., 2024) proposed to increase prompt complexity by letting the language model re-write the original prompts with increasing complexity. Other specific methods of creating complex synthetic data have been proposed, for example multi-hop question answering (Lupidi et al., 2024) or difficult reasoning questions (Yuan et al., 2025) both grounded on documents. Synthetic data has also been used to help train agents (Zhao et al., 2025; Zhou et al., 2025) and tool-use models (Mekala et al., 2024), and for rewriting pre-training data (Maini et al., 2024; Nguyen et al., 2025).
@@
Data selection is a critical component for post-training with synthetic data (and for data in general). Previously, LLM training was regarded as largely dependent on the size of available training data (Mishra et al., 2021; Wei et al., 2021; Wang et al., 2022b). More recent work has revealed that training on a smaller yet higher-quality curated set of prompts tends to be more effective in improving models’ both instruction following and reasoning capabilities (Zhou et al., 2024; Chen et al., 2024; Muennighoff et al., 2025; Ye et al., 2025). In addition to preprocessing techniques such as deduplication of similar prompts using similarity metrics such as ROUGE-L similarity score (Wang et al., 2022a) or clustering (Chen et al., 2023), as language models become more powerful, prompt curation can also be facilitated by using LLMs themselves as a quality judge. Recent work studies employing powerful language models to measure the complexity, diversity and quality of instructions (Lu et al., 2023; Chen et al., 2024; Touvron et al., 2023; Dubey et al., 2024; Li et al., 2023a). The success of RLHF for post-training (Stiennon et al., 2020; Rafailov et al., 2024) has attracted more attention to collecting large scale and high quality preference data. Most work involving preference optimization employs existing methods derived from pretraining and instruction-tuning (Touvron et al., 2023; Muennighoff et al., 2025), such as deduplication, clustering, quality classifiers or filtering heuristics. Rejecting Instruction Preferences (RIP) (Yu et al., 2025) is a recent method that gives strong performance by leveraging reward models based on the LLM responses to filter prompts. For verifiable reasoning tasks, Self-Consistency filtering (Prasad et al., 2024) has also been shown to be a high quality curation method by rejecting prompts where LLM solutions show no agreement, as the task is either incorrectly labeled or too difficult.
@@ 3 Chain-of-Thought(CoT)-Self-Instruct
(CoT)-Self-Instruct is an approach to generate high quality synthetic data for training using reasoning. We first assume access to a language model, and a small amount of high quality human-annotated seed data. We consider both verifiable reasoning domains, and non-verifiable general instruction following. Our approach involves two stages:
@@
    Synthetic Instruction Creation with Chain-of-Thought (CoT): given sample human-annotated seed instructions, we instruct the LLM to reason step by step to come up with instructions of similar complexity and domain.
@@
We then train LLMs using the generated high quality synthetic instructions. We describe each stage in turn.
@@ 3.1 Synthetic Instruction Creation via CoT
The process of CoT-Self-Instruct data creation starts with a small set of seed instructions as the instruction pool. Multiple instructions are sampled at random from the instruction pool, and then used to few-shot prompt a language model to generate a series of intermediate reasoning steps, followed by a new instruction. Unlike standard Self-Instruct (Wang et al., 2022a) which directly prompts the model to write new instructions given a list of seed instructions, each time we show the LLM an $N$-shot set of sample instructions, we first ask it to carefully analyze the given instructions, such as domain, complexity and purpose. After analyzing the seed instructions, reflecting on what makes them high quality prompts, the LLM is prompted to reason step by step to come up with a plan to generate a new self-contained instruction that is of similar quality and complexity as the given seed instructions, and ultimately to output the final synthetic instruction satisfying these requirements in a strict answer format.
@@
For reasoning tasks where there is a deterministic answer which we can compare against to generate verifiable rewards during training, we instruct the LLM to use reasoning to generate both an instruction and the verifiable target. The prompt we used for CoT-Self-Instruct on reasoning tasks is given in Figure 2.
@@
For tasks involving general instruction-following with open-ended responses, we direct the LLM to use reasoning to generate only the instruction, not the response itself. In these instances, later during training on this synthetic data we utilize a reward model to assess the responses, eliminating the need for a reference answer. The prompt we used for CoT-Self-Instruct on general instruction following tasks is given in Figure 3. Seed prompt pools for instruction-following typically include various different domains. When selecting few-shot prompts, for example combining prompts from storytelling and coding could result in unnatural synthetic prompts. To address this, we propose to first label seed prompts into categories, and then sample from the categories first.
Figure 2: CoT-Self-Instruct prompt generation template for verifiable reasoning tasks.
@@
Figure 3: CoT-Self-Instruct prompt generation template for general instruction following tasks.
@@ 3.2 Synthetic Instruction Curation
Even with the strongest language models, not all generated synthetic instructions are well-defined and answerable, or are effective in base model self-training. We therefore apply a curation step to select higher quality synthetic instructions from the pool of generated data for final post-training with RL.
@@
We propose Answer-Consistency to filter and retain only high quality data. Given the task instruction, we first instruct the LLM to generate $K$ responses and take the majority response. We then reject the data example and remove it from the training pool if the majority response does not match the target answer in the synthetic data example generated by CoT-Self-Instruct (i.e., via Figure 2). Because the target answer from CoT-Self-Instruct is generated with extensive reasoning steps during instruction writing and question-answering (Figure 2) this differs from the way a response is generated at inference time and might be more accurate. Hence, comparing if the two labels match gives an extra layer of filtering. We confirm in our experiments that Answer-Consistency is superior to standard Self-Consistency filtering (Prasad et al., 2024).
@@
For non-verifiable tasks, the Answer-Consistency method is not applicable. Instead, we employ the Rejecting Instruction Preferences (RIP) method as proposed by Yu et al. (2025). In this method, for a given task instruction, $K$ responses are generated, and each response is evaluated using a reward model (RM), resulting in a score for each response. The filtering process is then based on the distribution of these scores. In our setting we use the lowest score among these $K$ responses to represent the score of the synthetic prompt. We then filter the data by selecting only those prompts with the higher scores. Notably, this approach can also be applied to verifiable tasks, and we conduct experiments in that context as well.
@@ 3.3 Self-training with synthetic data
After generating the synthetic training data, we can conduct RL training on the set of generated instructions. We compare the performance of self-trained LLMs with models trained on human-annotated data and on seed instructions in reasoning and non-reasoning domains respectively. For verifiable reasoning tasks, we use GRPO (Shao et al., 2024), and for general instruction following we consider both offline DPO (Rafailov et al., 2024) and online DPO, which can perform much better, see e.g. Lanchantin et al. (2025).
@@ 4 Experimental Setup
We study the effectiveness of our synthetic prompt generation approach for reasoning and non-reasoning domains along the following two axes: synthetic prompt generation, and prompt curation.
4.1 Reasoning
@@
We use s1k (Muennighoff et al., 2025) reasoning instructions as our seed reasoning tasks. The s1k dataset consists of 1000 high-quality, diverse and difficult reasoning prompts. To conduct self-training with verifiable rewards we select a subset of s1k consisting of 893 verifiable reasoning instructions by filtering out theorem-proving questions and only keeping those that yield a scalar, single-valued, or simple closed-form answers that can be easily verified (such as 111, $A$, $False$, $\frac{2(n-1)(n-2)}{n(n+1)}$). We then use this subset as the seed instruction pool to generate more verifiable reasoning instructions.
Prompt Generation
The CoT-Self-Instruct template is given in Figure 2, and the baseline Self-Instruct method is given in Appendix Figure 6. To evaluate how CoT-Self-Instruct compares to baselines for generating verifiable reasoning tasks, we apply these methods to Qwen3-4B-Base models, Qwen3-4B model with Think mode and Qwen3-4B model with NoThink mode (Yang et al., 2025). We use temperature = 0.7 and top-p=0.8 for Qwen3-4B-Base and Qwen3-4B (NoThink mode), and temperature = 0.6 and top-p=0.95 for Qwen3-4B (Think mode).
@@
All our reasoning experiments use GRPO training initialized from Qwen3-4B-Base with reinforcement learning from rule-based verifiable rewards (RLVR). For hyperparameters, we use a cosine learning rate scheduler with a peak value of $1e-6$ and adopt the AdamW optimizer for the policy model. We set the number of training epochs to 40 with a batch size of 128. For rollouts, we sample 16 rollouts for each prompt with temperature = 0.6 and top-p=0.95, with a maximum length of 4096 tokens. All GRPO experiments are conducted with VeRL(Sheng et al., 2024) and Math-Verify  [^1] https://github.com/huggingface/Math-Verify as verifier.
@@
We construct targets for Self-Instruct and CoT-Self-Instruct when building the generated instructions. As a baseline, we also train on the original s1k prompts rather than synthetic data, in that case we use the public available DeepSeek R1 (Guo et al., 2025) thinking solution from simplescaling/s1K-1.1 to build targets. We also compare to training on OpenMathReasoning which consists of 10k prompts (Moshkov et al., 2025) with publicly available solutions by DeepSeek-R1 and QwQ-32B. We also explore some alternative ways of filtering data or constructing targets, explored as variations on our main experiments:
  • Self-Consistency filtering: generating responses $K$ times with random seeds and then selecting the majority-voted answer as the target or rejecting the example if the majority answer receives fewer votes than a given threshold (50% in our main experiments).
  • RIP filtering: we use the infly/INF-ORM-Llama3.1-70B (Minghao Yang, 2024) RM.
@@
We evaluate using Math500 (Hendrycks et al., 2021; Lightman et al., 2023), AIME 2024, AMC 23, and GPQA Diamond (Rein et al., 2024). We use temperature 0.6 and top-p 0.95 to generate predictions. For each problem we generate N = 16 solutions and report average accuracy.
4.2 Non-verifiable instruction following

### v2

# CoT-Self-Instruct: Building high-quality synthetic data for reasoning and non-reasoning tasks

@@ Abstract
We propose CoT-Self-Instruct, a synthetic data generation method that instructs LLMs to first reason and plan via Chain-of-Thought (CoT) based on given seed tasks, and then generate a new synthetic example of similar quality and complexity. This is followed by a filtering step to select high-quality data using automatic metrics, which are then used for LLM training. In verifiable reasoning, our synthetic data significantly outperforms existing training datasets, such as s1k and OpenMathReasoning, when evaluated on MATH500, AMC23, AIME24, and GPQA-Diamond. For non-verifiable instruction-following tasks, our method surpasses the performance of both human and standard Self-Instruct training data on the AlpacaEval 2.0 and Arena-Hard benchmarks.
@@ 1 Introduction
Synthetic data is artificially generated to replicate the characteristics and patterns of real-world data. One innovative approach to creating such data is the Self-Instruct method (Wang et al., 2022a), which utilizes LLMs themselves to generate instruction-following examples. This method begins by selecting a small set of seed instruction-following samples, which are then used to prompt LLMs to produce additional demonstrations in a similar format. Since then, a number of variants have been introduced that increase the complexity of queries (Liu et al., 2023; Zeng et al., 2024), maintain semantic diversity (Ding et al., 2023), scale the synthetic data (Yuan et al., 2023), and use these methods in self-improvement loops (Yuan et al., 2024). However, a significant challenge with these approaches is to ensure the quality and effectiveness of the generated data for language model training. Overall, generating high-quality synthetic data and optimizing its use for both reasoning and non-reasoning tasks still remains insufficiently understood.
@@
In this paper, we present Chain-of-Thought(CoT)-Self-Instruct, a method that both (i) uses reasoning to help create high-quality synthetic data; and (ii) self-filters the created data to only keep the highest quality ones; see Figure 1. Reasoning with CoT allows the model to analyze the given few-shot examples and plan the generation of challenging examples, while ensuring their logical validity. We show the efficacy of this approach for creating both verifiable reasoning data and non-verifiable instruction following tasks, where in both cases using CoT outperforms those generated without CoT. To curate high-quality verifiable data, we introduce Answer-Consistency, whereby we first create both a synthetic instruction and a target answer. Then we discard examples where this target answer does not match the majority vote solution of the LLM, with the assumption that those examples are either incorrectly labeled or too difficult. For non-verifiable data, we use the recent Rejecting Instruction Preferences (RIP) (Yu et al., 2025) method, which measures the quality of instructions based on the distribution of reward model scores assigned to sampled LLM responses. In both cases, filtering provides further gains. For reasoning tasks, CoT-Self-Instruct generated training data outperforms Self-Instruct and existing curated datasets such as s1k (Muennighoff et al., 2025) and OpenMathReasoning (Moshkov et al., 2025) when the trained LLM is evaluated on MATH500, AMC23, AIME24 and GPQA-Diamond. For non-verifiable tasks, it outperforms human data from WildChat (Zhao et al., 2024) and Self-Instruct synthetic data, whether filtered or not, on both the AlpacaEval 2 and ArenaHard benchmarks. Overall, models trained with CoT-Self-Instruct generated training data provided the best results among the methods we tested.
@@
Figure 1: CoT-Self-Instruct. Our method first prompts LLMs to reason and generate new instructions given seed instructions, followed by automatic curation of high-quality data using either Answer-Consistency for verifiable reasoning tasks, or RIP (Yu et al., 2025) for non-verifiable tasks.
@@ 2 Related Work
Synthetic data is produced using algorithms (Saxton et al., 2019), generative models (Borisov et al., 2022; Meng et al., 2022), or simulations (Vezhnevets et al., 2023), rather than being directly created by humans (Liu et al., 2024). It presents a promising solution for training models, particularly in scenarios where real-world data is scarce, expensive, or difficult to obtain. Self-Instruct (Wang et al., 2022a) first proposed a framework that prompts a language model with seed data as few-shot examples in order to generate new synthetic data. Such data has subsequently been used to self-train language models, e.g. in the Self-Rewarding framework (Yuan et al., 2024; Wu et al., 2024). Evol Instruct (Zeng et al., 2024) proposed to increase instruction complexity by letting the language model rewrite the original instructions with added complexity. Other specific methods of creating complex synthetic data have been proposed, such as multi-hop question answering (Lupidi et al., 2024) or difficult reasoning questions (Guo et al., 2025b; Yuan et al., 2025), both grounded on real documents. Synthetic data has also been used to help train agents (Zhao et al., 2025; Zhou et al., 2025) and tool-use models (Mekala et al., 2024), as well as for rewriting pre-training data (Maini et al., 2024; Nguyen et al., 2025).
@@
Data selection is a critical component for post-training with synthetic data (and for data in general). Previously, LLM training was regarded as largely dependent on the size of available training data (Mishra et al., 2021; Wei et al., 2021; Wang et al., 2022b). More recent work, however, has revealed that training on a smaller yet higher-quality curated set of instructions tends to be more effective in improving models’ both instruction following and reasoning capabilities (Zhou et al., 2024; Chen et al., 2024; Muennighoff et al., 2025; Ye et al., 2025). In addition to preprocessing techniques such as deduplication of similar instructions using similarity metrics such as ROUGE-L similarity score (Wang et al., 2022a) or clustering (Chen et al., 2023), as language models become more powerful, data curation can also be facilitated by using LLMs themselves as a quality judge. Recent work studies employing powerful language models to measure the complexity, diversity and quality of instructions (Lu et al., 2023; Chen et al., 2024; Touvron et al., 2023; Dubey et al., 2024; Li et al., 2023a). The success of RLHF for post-training (Stiennon et al., 2020; Rafailov et al., 2024) has attracted more attention to collecting large-scale and high-quality preference data. Most work involving preference optimization employs existing methods derived from pretraining and instruction-tuning (Touvron et al., 2023; Muennighoff et al., 2025), such as deduplication, clustering, quality classifiers or filtering heuristics. Rejecting Instruction Preferences (RIP) (Yu et al., 2025) is a recent effective method that leverages reward model scores assigned to LLM generated responses when filtering out low-quality instructions. For verifiable reasoning tasks, Self-Consistency filtering (Prasad et al., 2024) has also been shown to be a high-quality curation method by rejecting instructions where LLM solutions show low agreement, which suggests the task is either incorrectly labeled or too difficult.
@@ 3 Chain-of-Thought(CoT)-Self-Instruct
CoT-Self-Instruct is an approach to generate high-quality synthetic data for training using reasoning. We first assume access to a language model, and a small amount of high-quality human-annotated seed data. We consider both verifiable reasoning domains, and non-verifiable general instruction following. Our approach involves two stages:
@@
    Synthetic Instruction Creation with Chain-of-Thought (CoT): given sample human-annotated seed instructions, we instruct the LLM to reason step-by-step to come up with instructions of similar complexity and domain.
@@
We then train LLMs using the generated high-quality synthetic instructions. Below, we describe each stage in turn.
@@ 3.1 Synthetic Instruction Creation via CoT
The process of CoT-Self-Instruct data creation starts with a small set of seed instructions as the instruction pool. Multiple instructions are sampled at random from the instruction pool, and then used to few-shot prompt a language model to generate a series of intermediate reasoning steps, followed by a new instruction. Unlike standard Self-Instruct (Wang et al., 2022a), which prompts the model to directly write new instructions given a list of seed instructions, we first ask the model to carefully analyze the given instructions, such as their domain, complexity and purpose, and to reflect on what makes them high-quality instructions. After this analysis, the LLM then reasons step-by-step to come up with a plan to generate a new self-contained instruction that is of similar quality and complexity as the given seed instructions, and ultimately outputs the final synthetic instruction satisfying these requirements in a strict answer format.
@@
For reasoning tasks where there is a deterministic answer which we can compare against to generate verifiable rewards during training, we instruct the LLM to use reasoning to generate both an instruction and its verifiable answer at the same time. This allows the model to simultaneously solve the problem while creating it step-by-step, which can be easier than solving the final problem directly. The prompt we used for CoT-Self-Instruct on reasoning tasks is given in Figure 2, which has a strict formatting rule to allow easy separation of an instruction from its target answer.
@@
For tasks involving general instruction-following with open-ended responses, we direct the LLM to use reasoning to generate an instruction only, and not to include a specific response. In these instances, later during training on this synthetic data, we utilize a reward model to assess policy model responses, eliminating the need for a reference answer. The prompt we used for CoT-Self-Instruct on general instruction following tasks is given in Figure 3. Seed instruction pools for instruction-following typically include various different domains. When selecting few-shot samples, for example, combining instructions from storytelling and coding could result in unnatural synthetic instructions. To address this, we propose to first group seed instructions into categories by their domain. For each synthetically generated instruction, we first sample a category, and then select instructions from within that category as few-shot examples.
Figure 2: CoT-Self-Instruct instruction generation template for verifiable reasoning tasks.
@@
Figure 3: CoT-Self-Instruct instruction generation template for general instruction following tasks.
@@ 3.2 Synthetic Instruction Curation
Even with the strongest language models, not all generated synthetic instructions are well-defined and answerable, or are effective in base model self-training. We therefore apply a curation step to select higher-quality synthetic instructions from the pool of generated data for final post-training with RL.
@@
We propose Answer-Consistency to filter and retain only high-quality data. Given the task instruction, we first instruct the LLM to generate $K$ responses and take the majority vote over the final answers. We then reject the data example and remove it from the training pool if the majority vote does not match the target answer included in the synthetic data example generated by CoT-Self-Instruct (i.e., via Figure 2). Because the target answer from CoT-Self-Instruct is generated with extensive reasoning steps during instruction writing and question-answering (Figure 2), it has access to more information (e.g. the step-by-step creation process of a problem) compared to an LLM answer generated at inference time given only the problem. Hence, comparing if the answers match gives an extra layer of filtering. We confirm in our experiments that Answer-Consistency is superior to standard Self-Consistency filtering (Prasad et al., 2024).
@@
For non-verifiable tasks, the Answer-Consistency method is not applicable as we do not generate target responses when creating synthetic examples for open-ended questions. Instead, we employ ideas from the Rejecting Instruction Preferences (RIP) method as proposed by Yu et al. (2025). In this method, for a given task instruction, $K$ responses are generated, and each response is evaluated using a reward model (RM), resulting in a rating for each response. The filtering process is then based on the distribution of these ratings. We use the lowest rating among these $K$ responses to represent the overall score of a given synthetic instruction. While Yu et al. (2025) set a global threshold for filtering over these scores, we notice a topic distribution shift if we perform such global filtering, as each topic category has different score distributions. Therefore, we slightly modify RIP by sampling multiple instructions from each few-shot prompt, then selecting the one with the highest score. Notably, the RIP approach can also be applied to verifiable tasks, and we conduct experiments in that context as well.
@@ 3.3 Self-training with Synthetic Data
After generating the synthetic training data, we can then use them to conduct RL training in order to improve the downstream performance of an LLM. We compare the performance of such self-trained LLMs with models trained on human-annotated data and on curated seed instructions in reasoning and general instruction following domains respectively. For verifiable reasoning tasks, we train with GRPO (Shao et al., 2024), and for general instruction following we consider both offline DPO (Rafailov et al., 2024) and online DPO, which can perform much better, see e.g. Lanchantin et al. (2025).
@@ 4 Experimental Setup
We study the effectiveness of our synthetic data generation approach for reasoning and general instruction following domains along the following two axes: synthetic instruction generation, and instruction curation.
4.1 Verifiable Reasoning Tasks
@@
We use the s1k reasoning instructions from (Muennighoff et al., 2025) as our seed tasks. The s1k dataset contains 1,000 high-quality, diverse, and challenging reasoning questions, which are on par with those found in the original 59,000 sample dataset according to their paper. To conduct self-training with verifiable rewards, we select a subset of s1k consisting of 893 verifiable reasoning instructions by filtering out theorem-proving questions and only keeping those that yield a scalar, single-valued, or simple closed-form answers that can be easily verified (such as $1$, AA, FalseFalse, $\frac{2(n-1)(n-2)}{n(n+1)}$). We then use this subset as the seed instruction pool to generate more verifiable reasoning instructions.
Instruction generation
The CoT-Self-Instruct template is given in Figure 2. To evaluate how CoT-Self-Instruct compares to baselines for generating verifiable reasoning tasks, we apply these methods to the following models: Qwen3-4B-Base, Qwen3-4B with Think mode and Qwen3-4B with NoThink mode (Yang et al., 2025). We generate up to 10,000 instructions using temperature=0.7 and top-p=0.8 for Qwen3-4B-Base and Qwen3-4B (NoThink mode), and temperature=0.6 and top-p=0.95 for Qwen3-4B (Think mode).
@@
All our reasoning experiments use GRPO training initialized from Qwen3-4B-Base with reinforcement learning from rule-based verifiable rewards (RLVR). For hyperparameters, we use a cosine learning rate scheduler with a peak value of 1e61e-6 and adopt the AdamW optimizer for the policy model. We set the number of training epochs to 40 with a batch size of 128. For rollouts, we sample 16 rollouts for each prompt with temperature=0.6 and top-p=0.95, with a maximum length of 4096 tokens. All GRPO experiments are conducted with VeRL(Sheng et al., 2024) and Math-Verify  [^1] https://github.com/huggingface/Math-Verify as a verifier.
@@
We compare our method CoT-Self-Instruct with standard Self-Instruct with no CoT for generating either instruction or target answer (template given in Appendix Figure 7). We also compare to using no CoT for generating the instruction, but using CoT to generate the target (template given in Appendix Figure 6). These variations test the importance of using CoT for both instruction and target when generating synthetic training examples. As further baselines, we also train on the original s1k instructions rather than synthetic data, in that case we use the publicly available DeepSeek R1 (Guo et al., 2025a) thinking solution from simplescaling/s1K-1.1 to build targets. We also compare to training on OpenMathReasoning which consists of 10k instructions (Moshkov et al., 2025) with publicly available solutions by DeepSeek-R1 and QwQ-32B. We also explore some alternative ways of filtering data or constructing target labels, explored as variations on our main experiments:
  • Self-Consistency filtering: generate $K$ responses with random seeds and then select the majority-voted answer as the target or reject the example if the majority answer receives fewer votes than a given threshold (50% in our main experiments).
  • RIP filtering: we use the infly/INF-ORM-Llama3.1-70B (Minghao Yang, 2024) reward model.
@@
We evaluate our trained models on math tasks using MATH500 (Hendrycks et al., 2021; Lightman et al., 2023), AIME 2024 and AMC 23. We also evaluate on GPQA Diamond (Rein et al., 2024), which consists of challenging science questions. We use temperature=0.6 and top-p=0.95 to generate predictions. For each problem, we generate $N=16$ solutions and report the average accuracy.
4.2 General Instruction Following